A block slides down an incline of 30° with an acc. g/4.find the kinetic friction coefficient.

Grade:12

5 Answers

vishal lakha
12 years ago

very simple funda...tan theta =coeff. of kinetic friction

tan 30=1/underroot 3.

answer.

-vishal lakha

punjab.

PHANI TEJA
12 years ago

tan theta = the coefficient of kinetic friction.

tan30 =1/3 under root =coeficient of kietic friction



varadharajan sridharan
12 years ago

tan 30 = (1/3)^(1/2)

ranjeet kumar
12 years ago

what is water drops are falling from a tap at regular time.when frist drop tou

ranjeet kumar
12 years ago

A block slides down an incline of 30° with an acc. g/4.find the kinetic friction coefficient.